### v1

@@ 3.4 Evaluation metrics
To conduct a quantitative evaluation, we selected four authoritative metrics: the Coefficient of Determination ($\text{R}^{2}$), Root Mean Square Error (RMSE), Peak Signal-to-Noise Ratio (PSNR)Hore and Ziou (2010), and the Universal Image Quality Index (UIQI)Wang and Bovik (2002).
Among these, the Coefficient of Determination ($\text{R}^{2}$)is utilized to quantify the extent to which the generated data explains the variability of the reference NPP-VIIRS NTL data. It is calculated as follows:
$\text{R}^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$ (3)
where $y_{i}$ is the value of the $i$-th pixel in the reference image, $\hat{y}_{i}$ is the value of the corresponding pixel in the generated image, $\bar{y}$ is the mean value of all pixels in the reference image, and $n$ is the total number of pixels. A value closer to 1 indicates a better goodness of fit for the model, signifying that the generated data aligns well with the distribution characteristics of the real data.
The Root Mean Square Error (RMSE) assesses the numerical precision of the data by measuring the average deviation between predicted and actual values. The formula is as follows:
$\text{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}$ (4)
where $y_{i}$ represents the value of the $i$-th pixel in the reference image, $\hat{y}_{i}$ is the value of the corresponding pixel in the generated image, and $n$ is the total number of pixels. A lower RMSE value indicates higher precision of the generated data.
The Peak Signal-to-Noise Ratio (PSNR) assesses the reconstruction quality of an image by quantifying the ratio between the maximum possible signal power and the power of corrupting noise. It is calculated as:
$\text{PSNR}=10\cdot\log_{10}\left(\frac{MAX^{2}}{MSE}\right)$ (5)
where MAX is the maximum possible pixel value of the image, and MSE is the mean squared error defined as $\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}$, with $y_{i}$ representing the value of the $i$-th pixel in the reference image, $\hat{y}_{i}$ the value of the corresponding pixel in the generated image, and $n$ the total number of pixels. PSNR is expressed in decibels (dB). A higher PSNR value indicates a lower MSE, signifying that the quality of the generated image is higher and its distortion relative to the reference image is lower.
Finally, the Universal Image Quality Index (UIQI) provides a comprehensive assessment of image quality by integrating three factors: loss of correlation, luminance distortion, and contrast distortion. It is computed as follows:
$\text{UIQI}=\frac{4\sigma_{y\hat{y}}\bar{y}\bar{\hat{y}}}{(\sigma_{y}^{2}+\sigma_{\hat{y}}^{2})(\bar{y}^{2}+\bar{\hat{y}}^{2})}$ (6)
where $\bar{y}$ and $\bar{\hat{y}}$ are the mean values of the reference image $y$ and the generated image $\hat{y}$, respectively; $\sigma_{y}^{2}$ and $\sigma_{\hat{y}}^{2}$ are their respective variances; and $\sigma_{y\hat{y}}$ is the covariance between the reference and generated images. The UIQI index has a value range of [1,1][-1,1][ - 1 , 1 ], where a value closer to 1 signifies higher quality in the generated image and greater similarity to the reference image.
3.5 Implementation Details
All models were implemented in Python using PyTorch 2.6.0 + CUDA 12.6. Training was carried out on a single NVIDIA RTX 3090 Ti GPU (24 GB VRAM); results were cross-checked on an NVIDIA L40S to ensure hardware-independence. [^1] Reproducing the experiments on either GPU yielded numerically consistent validation metrics (differences $<10^{-4}$ in loss after convergence).
Data sampling:
[…]
Pre-processing:
All NTL and Landsat inputs were resampled to 250 m; the GAIA mask was ingested at 50 m and downsampled on-the-fly within the Dual-Feature Refiner (DFR) to match the current scale of each attention gate. NPP-VIIRS targets were natural-log transformed during training and exponentiated at inference time. Input channels were min–max normalized to [0,1][0,1][ 0 , 1 ] per band across the training set; the same statistics were applied to validation/test splits.
Network configuration:
The backbone encoder follows a ResNet-50 initialization (ImageNet pre-trained weights). After the final encoder stage, feature maps are projected to a channel pyramid of {512, 256, 128, 64, 32} to feed the symmetric decoder through skip connections. The decoder incorporates the proposed HFD at each upsampling level (Section 3.2). The DFR is attached post-backbone and trained to refine high-frequency structural cues using the Cross-Resolution Local Attention guided by the impervious-surface mask (Section 3.3).
Optimization:
[…]
Model selection and inference:
Checkpoint selection was based on the lowest validation RMSE (computed in log space). The selected weights were used to generate annual EVAL products (starting from 1986) following the two-phase sensor strategy in Section 2. Outputs were written at 500 m resolution to balance storage and downstream analysis; per-pixel exponentiation restored radiance units.
4 Results
@@ 4.1 Accuracy assessment
LongNTLChen et al. (2020) 0.5961 1.4483 43.5045 0.6796
SVNLChen et al. (2024) 0.6857 1.2776 44.5936 0.7726
@@
4.2 Validation of early-period time-series
@@
We generated the EVAL dataset in two distinct phases. For the 1992–2013 period, we used the Harmonized Global Nighttime Light dataset, whose reliability has been established in prior tests. In contrast, for the previous period from 1986 to 1991, we used the PANDA-China dataset as input, which required additional validation to confirm its robustness. To perform this validation, we designed a proxy evaluation. We assessed our model’s performance from 2014 to 2018, a period where the framework also relies on PANDA-China inputs. We then compared these generated results against actual NPP-VIIRS data from the same years. This approach allows us to infer the model’s likely performance during 1986 to 1991.
Table. 3 presents the results for this proxy-period evaluation. EVAL’s accuracy remained high and stable from 2014 to 2016. During these years, the pixel-scale Coefficient of Determination ($\text{R}^{2}$) stayed around 0.7, peaking in 2014 ($\text{R}^{2}$ =0.7743), while the RMSE was consistently below 1.4. At the city scale, the $\text{R}^{2}$ values were even stronger, generally exceeding 0.9. However, the model’s accuracy declined from 2017 to 2018. At the pixel scale, the $\text{R}^{2}$ dropped sharply to below 0.6, and the RMSE rose above 1.5. This performance drop aligns with the known behavior of the PANDA-China dataset. Its predictions rely on time-series models, and their accuracy decreases as the temporal distance from the base years increases. The reduced performance of EVAL after 2017 likely reflects this inherited limitation.
Based on these findings, we can make clear recommendations for using the early-period data. We infer that the EVAL dataset provides reliable performance at the grid scale from 1988 to 1991. For the earliest years of 1986 and 1987, we recommend using the dataset at the city scale for the most robust analysis.
Years $\text{R}^{2}_{\text{pixel}}$ $RMSE_{\text{pixel}}$ PSNR UIQI $\text{R}^{2}_{\text{city}}$ $\text{RMSE}_{\text{city}}$ 2014 0.7743 1.1347 44.9635 0.8718 0.9606 1910.4 2015 0.7360 1.2386 43.0582 0.8412 0.9510 2171.8 2016 0.6978 1.3937 45.1749 0.8043 0.9275 2692.9 2017 0.6043 1.7189 43.4423 0.7177 0.8851 3738.3 2018 0.6135 1.8640 42.7856 0.6668 0.8045 5354.7
Table 3: Quantitative evaluation of early-period time-series.
4.3 Validation of time-series consistency
@@
To assess the temporal consistency of the EVAL, this study conducted a comparative analysis with several established long-term NTL datasets in Figure. 6. We statistically analyzed and compared the trends in the total annual NTL values for China across these datasets to evaluate EVAL’s overall coherence and relative performance in capturing long-term light dynamics. The analysis reveals that the time series of EVAL is only marginally shorter than that of the PANDA-China dataset, which is based on DMSP-OLS NTL data. When compared to other NPP-VIIRS-like NTL datasets, EVAL demonstrates a significant advantage in temporal coverage, offering unique support for high-precision, long-term studies. Furthermore, EVAL exhibits a high degree of temporal consistency across its entire time range, characterized by a stable growth trend. From 1992 to 2013, the trend in EVAL closely aligns with that of the DMSP-OLS NTL dataset. Additionally, EVAL’s trend of variation is similar to that of existing VIIRS-like NTL products, but its values are significantly higher. This results in a stronger consistency with the official NPP-VIIRS NTL data, which in turn avoids the abrupt increase seen in LongNTL and SVNL during the 2012–2013 transition to the NPP-VIIRS record.
@@
Figure 6: Comparison of Annual NTL Sums from different datasets. EVAL, SVNL, LongNTL, and NPP-VIIRS use the left y-axis (Total NTL in $W/cm^{2}\cdot sr$), while PANDA-China and DMSP-OLS use the right y-axis (Total NTL in Digital Number).
@@
4.4 Socioeconomic correlation analysis of NTL products
Correlation analysis between NTL data and socioeconomic indicators is a vital method for evaluating its quality and utility. A strong correlation indicates that the data accurately reflects socioeconomic activities, providing substantial value for research in fields like economic monitoring and regional development. We therefore compared VIIRS-like products against two key socioeconomic indicators: Gross Domestic Product (GDP) and Resident Population (POP) in Table. 4. We selected the 2000–2012 period for this analysis, as it was the overlapping timeframe where all three datasets were available. The statistical results reveal that EVAL consistently outperformed the other products. Specifically, its correlation coefficient with GDP was 0.9681, compared to 0.8622 for LongNTL and 0.8997 for SVNL. Similarly, its correlation with POP was 0.9402, again surpassing LongNTL (0.8846) and SVNL (0.9270). On average, EVAL’s correlation coefficient reached 0.9542, representing an improvement of 9.3% over LongNTL (0.8734) and 4.5% over SVNL (0.9134). This strong performance underscores EVAL’s superior capacity to capture the intensity of human activity and economic development, making it a more reliable and applicable tool for socioeconomic research.
Table 4: Correlation Analysis of NTL Products with socioeconomic parameters.
Socioeconomic Indicators LongNTL SVNL EVAL
Gross Domestic Product 0.8622 0.8997 0.9681
Resident population 0.8846 0.9271 0.9402
Average 0.8734 0.9134 0.9542
5 Discussion
5.1 Ablation study
@@
5.2 Limitations
Despite the promising results, this study has several limitations that should be acknowledged. First, the accuracy and reliability of the EVAL dataset during the early period of its time series warrant further evaluation and enhancement. In this period, EVAL relies on the PANDA-China dataset for its input features and may therefore inherit its inherent predictive biases and uncertainties, which could constrain the accuracy of the generated results. Furthermore, the EVAL dataset has not yet undergone a rigorous time-series consistency calibration at the pixel level. This may cause the nighttime light intensity values of individual pixels to exhibit abrupt changes or irregular fluctuations over time, potentially interfering with the findings of fine-scale longitudinal studies. Finally, the present study has only produced nighttime light data covering the geographic extent of China. This scope limits the application potential of the EVAL dataset for studies conducted on larger regional or global scales.
6 Conclusion
This paper introduces a novel framework for reconstructing NPP-VIIRS-like NTL data from multi-source data. Based on this framework, we produce a new NPP-VIIRS-like NTL product for China, named EVAL, which spans the period from 1986 to the present. Our proposed framework is composed of a construction stage and a refinement stage. To address the underestimation and omission phenomena prevalent in existing products, we incorporate the Structure Residual Fusion (SRF) module and the Multi-Scale Aggregator (MA) into the Hierarchical Fusion Decoder (HFD), and design a Dual Feature Refiner (DFR) featuring a Cross-Resolution Local Attention to enhance the prediction of localized fine-grained features. The efficacy of our product is demonstrated through a comparative analysis against two state-of-the-art products using four distinct evaluation metrics. The experimental results verify the comprehensive superiority of EVAL, while a series of ablation studies further confirm that our designed modules effectively mitigate the issues of underestimation and omission.
Nevertheless, our framework and the resulting product have several limitations. For instance, artifacts are present in some of the generated images. The accuracy of EVAL between 1986 and 1991 warrants further improvement, and the dataset has not yet undergone a rigorous pixel-level time-series consistency calibration. In our future work, we plan to address these limitations sequentially. We also intend to expand the spatial coverage of our product to generate a global version of the EVAL dataset.
In conclusion, our framework offers a new approach for the reconstruction of NPP-VIIRS-like NTL data. Moreover, as the longest available time-series NPP-VIIRS-like NTL product to date, EVAL can provide powerful support for a range of applications, including human activity monitoring, urbanization process assessment, and sustainable development goals.
7 Data available
The complete EVAL dataset, covering the full period from 1986 to 2024, is publicly available at https://doi.org/10.11888/HumanNat.tpdc.302930. The EVAL dataset is a composite product designed to provide a continuous, long-term time series. It consists of two distinct components corresponding to different periods:
Reconstructed Period (1986–2013): This part of the dataset contains the annual, 500m resolution VIIRS-like NTL data generated by our proposed reconstruction framework. Processed Official Period (2012–2024): This part consists of the official annual mean NPP-VIIRS DNB data, which we have processed by capping high-value pixels to ensure consistency. The overlapping years of 2012 and 2013 serve a critical purpose: they allow for a direct comparison between our reconstructed data and the official data, enabling a robust evaluation and validation of our model’s performance.
[…]

### v2

@@ Methods
To conduct a quantitative evaluation, we selected four authoritative metrics: the Coefficient of Determination ($\text{R}^{2}$), Root Mean Square Error (RMSE), Peak Signal-to-Noise Ratio (PSNR)(hore2010image, ), and the Universal Image Quality Index (UIQI)(wang2002universal, ).
Implementation details
The model was implemented in Python using PyTorch and trained on a single NVIDIA RTX 3090 Ti GPU. It utilized 20,000 512x512 pixel image patches from a 2013 China-wide dataset, containing co-registered night-time light, Landsat, and impervious surface data. To address the data imbalance from large dark areas, these patches were selected only if at least 1% of their pixels had non-zero values. These patches were split into 80/20 for training and validation. Input data was resampled to 250m, while the network architecture featured a ResNet-50 backbone, a symmetric decoder with the proposed HFD module, and a Dual-Feature Refiner (DFR) to enhance structural details. The optimization was a two-stage process: the main network was trained for 60 epochs using the Adam optimizer with a learning rate of 1e-4, followed by an additional 10 epochs of training for only the DFR module using L1 loss and a lower learning rate of 1e-5. To generate the final nationwide imagery, we used a sliding window approach with a 6-pixel overlap between adjacent patches. The average value in these overlapping regions was used as the final prediction to ensure a seamless and consistent reconstruction.
Data Records
The complete EVAL dataset, covering the full period from 1986 to 2024, is publicly available at the National Tibetan Plateau Data Center (TPDC) (https://doi.org/10.11888/HumanNat.tpdc.302930)tian2025extended . The EVAL dataset is a composite product designed to provide a continuous, long-term time series. It consists of two distinct components corresponding to different periods:
Reconstructed Period (1986-2013): This part of the dataset contains the annual, 500m resolution VIIRS-like NTL data generated by our proposed reconstruction framework. To produce this, we used two different DMSP-OLS inputs: for the 1992-2013 period, we used Li’s NTL; for the 1986-1991 period, we used PANDA-China. The reliability of the data from this earlier 1986-1991 period is specifically addressed in the ”Validation of early-period time-series” section.
Processed Official Period (2012-2024): This part consists of the official annual mean NPP-VIIRS DNB data, which we have processed by capping high-value pixels to ensure consistency.
The overlapping years of 2012 and 2013 serve a critical purpose: they allow for a direct comparison between our reconstructed data and the official data, enabling a robust evaluation and validation of our model’s performance.
Technical Validation
@@
LongNTL(chen2020extended, ) 0.5961 1.4483 43.5045 0.6796
SVNL(chen2024global, ) 0.6857 1.2776 44.5936 0.7726
@@
Validation of early-period time-series
@@
We generated the EVAL dataset in two distinct phases. For the 1992–2013 period, we used the Li’s Nighttime Light dataset as input, whose reliability has been established in prior tests. In contrast, for the previous period from 1986 to 1991, we used the PANDA-China dataset as input, which required additional validation to confirm its robustness.
To perform this validation, we designed a spatial stability evaluation using GDP and Resident Population (POP) data as proxies for human activity. The core assumption is that the spatial pattern of the relationship between NTL and these proxies should be very stable over time. Significant data quality issues, like sensor noise or calibration errors, would noticeably disrupt this stable spatial pattern.
Our method proceeds in two steps.
Normalization To remove the effect of overall economic growth, we compute a normalized statistic for each province $i$ in each year $j$:
$n_{ij}=\frac{\mathrm{NTL}_{ij}}{K_{ij}},$ (3)
where $\mathrm{NTL}_{ij}$ is the nighttime light intensity for province $i$ in year $j$, and $K_{ij}$ is an external proxy for the economic scale of that province in that year. Thus, $n_{ij}$ represents nighttime light intensity per unit of the proxy. We arrange all $n_{ij}$ values into a matrix $N=[n_{ij}]$, where each column
$N_{j}=(n_{1j},n_{2j},\ldots,n_{Ij})^{\top}$
is the spatial distribution across all $I$ provinces in year $j$.
Spatial consistency check For each year $j$, we quantify how similar its spatial pattern $N_{j}$ is to that of a reliable baseline year $N_{y}$ by computing the coefficient of determination ($R^{2}$) between $N_{j}$ and $N_{y}$. A high $R^{2}$ indicates that year $j$ reproduces the benchmark’s spatial structure well, suggesting high data quality with low noise. Conversely, a low $R^{2}$ indicates a substantial deviation from the benchmark pattern, suggesting lower data quality.
The quantitative results of this spatial stability evaluation are presented in Table 3. We first establish a benchmark using the original NPP-VIIRS satellite record (2013–2018). As these data show, the spatial relationship between modern NTL and the proxies is highly stable, with coefficients of determination ($R^{2}$) consistently high for both GDP ($R^{2}$ range: 0.930–0.983) and POP ($R^{2}$ range: 0.932–0.988). This stable range provides a standard for assessing the quantity of our reconstructed data.
Table 3: Quantitative evaluation of spatial stability. EVAL uses 1992 as the baseline year, while NPP-VIIRS uses 2012 as the baseline year.
EVAL NPP-VIIRS
YEAR GDP POP YEAR GDP POP YEAR GDP POP
1991 0.9609 0.9869 1993 0.9679 0.9881 2013 0.9827 0.9883
1990 0.9486 0.9911 1994 0.9417 0.9718 2014 0.9633 0.9625
1989 0.9385 0.9886 1995 0.9508 0.9683 2015 0.9702 0.9610
1988 0.9212 0.9782 1996 0.8736 0.9506 2016 0.9490 0.9551
1987 0.9103 0.9878 1997 0.8490 0.9322 2017 0.9477 0.9552
1986 0.8732 0.9860 1998 0.8523 0.9594 2018 0.9297 0.9323
Against this benchmark, our analysis of the EVAL dataset yields two key insights. First, EVAL demonstrates high internal consistency. The stability metrics for the 1986–1991 period (derived from PANDA-China) are shown to be highly comparable to those from the 1992–1998 period (derived from Li’s NTL). As seen in Table 3, the $R^{2}$ ranges for both GDP and POP are commensurate across the two periods. This indicates that the data source transition in 1992 did not introduce a systemic shock to the dataset’s reliability. More critically, the stability of the reconstructed 1986–1991 EVAL data is found to be directly comparable to the NPP-VIIRS standard. The $R^{2}$ metrics for EVAL in this earliest period are largely on par with the stability range observed in the actual NPP-VIIRS record. This result provides strong quantitative evidence for the consistency and reliability of the EVAL dataset, validating the quality of the data generated for the contested 1986–1991 period.
Validation of time-series consistency
@@
To assess the temporal consistency of the EVAL, this study conducted a comparative analysis with several established long-term NTL datasets in Figure. 6. We statistically analyzed and compared the trends in the total annual NTL values for China across these datasets to evaluate EVAL’s overall coherence and relative performance in capturing long-term light dynamics. The analysis reveals that the time series of EVAL is only marginally shorter than that of the PANDA-China dataset, which is based on DMSP-OLS NTL data. When compared to other NPP-VIIRS-like NTL datasets, EVAL demonstrates a significant advantage in temporal coverage, offering unique support for high-precision, long-term studies. Furthermore, EVAL exhibits a high degree of temporal consistency across its entire time range, characterized by a stable growth trend. From 1992 to 2013, the trend in EVAL closely aligns with that of the corrected DMSP-OLS NTL datasets. Additionally, EVAL’s trend of variation is similar to that of existing VIIRS-like NTL products, but its values are significantly higher. This results in a stronger consistency with the official NPP-VIIRS NTL data, which in turn avoids the abrupt increase seen in LongNTL and SVNL during the 2012–2013 transition to the NPP-VIIRS record.
@@
Figure 6: Comparison of Annual NTL Sums from different datasets. EVAL, SVNL, LongNTL, and NPP-VIIRS use the left y-axis (Total NTL in $nW/cm^{2}\cdot sr$), while PANDA-China and DMSP-OLS use the right y-axis (Total NTL in Digital Number).
@@
Socioeconomic correlation analysis
Correlation analysis between NTL data and socioeconomic indicators(SEI) is a vital method for evaluating its quality and utility. A strong correlation indicates that the data accurately reflects socioeconomic activities, providing substantial value for research in fields like economic monitoring and regional development. We therefore compared VIIRS-like products against two key socioeconomic indicators: Gross Domestic Product (GDP) and Resident Population (POP) in Table. 4. We selected the 2000–2012 period for this analysis, the overlapping timeframe where all seven datasets were available. Among the evaluated products, DMSP-OLS, as the original NTL data, serves as a baseline for performance, showing an average correlation of 0.9356 with socioeconomic indicators. Li’s NTL demonstrates an improvement upon this baseline through fine-tuning, achieving an average correlation of 0.9467.
Building upon Li’s NTL, the reconstructed EVAL dataset demonstrates a comprehensive enhancement in performance. It surpasses its predecessor across every metric, registering superior correlation coefficients for GDP (0.9681), POP (0.9402), and the overall average (0.9542). This result confirms the effectiveness of our proposed reconstruction framework.
Furthermore, EVAL’s superior performance extends to a broader comparison with other contemporary 500-meter resolution NTL products. The analysis reveals that EVAL consistently leads this category. Its correlation coefficients for both GDP and POP are the highest among all 500m products, outperforming Tongji-NTL (0.8649, 0.8352), LongNTL (0.8622, 0.8846), and SVNL (0.8997, 0.9271). This consistent leadership across all metrics underscores EVAL’s position as the most accurate and reliable 500m resolution NTL product in this study for socioeconomic applications.
Table 4: Correlation Analysis of NTL Products with Socioeconomic Indicators.
SEI DMSP-OLS Li’s NTL Tongji-NTL LongNTL SVNL EVAL
GDP 0.9475 0.9560 0.8649 0.8622 0.8997 0.9681
POP 0.9236 0.9374 0.8352 0.8846 0.9271 0.9402
AVG 0.9356 0.9467 0.8506 0.8734 0.9134 0.9542
We further leverage EVAL’s unique long-term nature to conduct a historical correlation analysis spanning from 1986. This segmented analysis, shown in Table 5, reveals how the relationship between NTL and socioeconomic activity has evolved over nearly three decades.
The correlation was already strong in the initial 1986-1991 period, showing an average value of 0.8943. This relationship then intensified significantly during the 1992-1999 phase of rapid economic growth, where the average correlation reached a peak of 0.9632. This high level of correlation was maintained through the 2000-2012 period (average of 0.9542), reinforcing the findings from our previous comparison.
Overall, for the entire 25-year span from 1986 to 2012, the EVAL dataset maintained a remarkable average correlation of 0.9216 with socioeconomic indicators and 0.9673 with GDP. This confirms not only EVAL’s superior performance against other products but also its robustness and reliability for long-term historical studies.
Table 5: Correlation Coefficients between the EVAL Dataset and Socioeconomic Indicators, 1986–2012
SEI 1986-1991 1992-1999 2000-2012 1986-2012
GDP 0.9003 0.9791 0.9681 0.9673
POP 0.8882 0.9472 0.9402 0.8758
AVG 0.8943 0.9632 0.9542 0.9216
Ablation study
@@
Data availability
The Extended VIIRS-like Artificial Nighttime Light (EVAL) dataset generated in this study is available at the National Tibetan Plateau Data Center (TPDC) (https://doi.org/10.11888/HumanNat.tpdc.302930tian2025extended ).
Code availability
The programs used to generate all the results were Python 3.12. The code and scripts used for training, testing, and predicting the NTL data are available in the open GitHub repository “https://github.com/Rookie1miao/EVAL_framework”.
Acknowledgements
This study is supported by National Key Research and Development Program of China (2022YFB3903703).
Author contributions
Yihe Tian: Conceptualization, Methodology, Software, Validation, Formal analysis, Visualization, Writing Kwan Man Cheng: Conceptualization, Methodology, Resources, Validation, Formal analysis Zhengbo Zhang: Resources, Validation Tao Zhang: Review & Editing Zhehao Ren: Review & Editing Junning Feng: Review & Editing Suju Li: Review & Editing Dongmei Yan: Review & Editing Bing Xu: Review, Editing, Formal analysis, Funding acquisition, Supervision.
Competing interests
The authors declare no competing interests.